rvc (rvc):

The acceleration of the train starting from the rest is a= 8/(v^2+1). Find a. v=? when x= 20m b. x=? when v=64.8km/hr

OpenStudy (michele_laino):

question a) the velocity is given by the subsequent formula: \[\Large v = at\]

rvc (rvc):

is it the differential equation?

OpenStudy (michele_laino):

please wait

OpenStudy (michele_laino):

yes! I think that we have to write a differential equation

rvc (rvc):

yes but im confused which to use

OpenStudy (michele_laino):

by definition of acceleration, we can write this: \[\Large a = \frac{{dv}}{{dt}} = \frac{8}{{{v^2} + 1}}\]

rvc (rvc):

brb

OpenStudy (michele_laino):

the space x(t) traveled is given by this differential equation: \[\Large \frac{{dx}}{{dt}} = v\left( t \right)\]

OpenStudy (michele_laino):

we have to integrate that ODE, so we have: \[\Large x\left( t \right) = \int {v\left( t \right)dt} \] then we have to know the shape of the function: \[\Large v = v\left( t \right)\]

rvc (rvc):

the question doesn't tell us about the shape

OpenStudy (michele_laino):

yes I know, we have to compute it, nevertheless it is not simple, since after a simple integration, we get: \[\Large \frac{{{v^3}}}{3} + v = 8t\] as you can easily check

OpenStudy (michele_laino):

sincerely I don't know how to overcome this difficulty

rvc (rvc):

we don't have t in the equation

OpenStudy (michele_laino):

please wait I ask to another helper

OpenStudy (michele_laino):

@mathstudent55 please help

OpenStudy (michele_laino):

@e.mccormick please help

rvc (rvc):

I think we have these formulas: \[\Large \rm a=v \frac{ dv }{ dx }\] \[\Large \rm v=\frac{ dx }{ dt }\] \[\Large \rm a=\frac{ dv }{ dt }\]

OpenStudy (michele_laino):

yes! nice idea!

OpenStudy (haseeb96):

it is very easy I think @Michele_Laino can figure out your this problem. I have to go on my roof. Because there is some huge storm coming.

OpenStudy (michele_laino):

we have: \[\Large \frac{{dx}}{{dv}} = \frac{v}{a} = \frac{{v\left( {{v^2} + 1} \right)}}{8}\]

rvc (rvc):

Okay @Haseeb96 Thanks bro :)

OpenStudy (michele_laino):

thanks @Haseeb96 actually @rvc helped me with her nice idea

OpenStudy (michele_laino):

ok! we have only to integrate that last ODE, so we can write this: \[\Large x\left( v \right) = \frac{1}{8}\left( {\frac{{{v^4}}}{4} + \frac{{{v^2}}}{2}} \right)\]

rvc (rvc):

wait how dx/dv? can we write like that way?

OpenStudy (michele_laino):

since: \[\Large \frac{{dx}}{{dv}} = \frac{1}{{\frac{{dv}}{{dx}}}}\]

rvc (rvc):

oh okay

OpenStudy (michele_laino):

after a substitution, we gave to solve a qudratic equation for v^2

OpenStudy (michele_laino):

we have*

OpenStudy (michele_laino):

for part b) we can use the same formula: \[\Large x\left( v \right) = \frac{1}{8}\left( {\frac{{{v^4}}}{4} + \frac{{{v^2}}}{2}} \right)\]

OpenStudy (michele_laino):

since the motion is the same. We have only to substitute v=64.8 at the right side

rvc (rvc):

and the first condition?

OpenStudy (michele_laino):

do you mean part a) ?

rvc (rvc):

yep

OpenStudy (michele_laino):

we have to substitute your data, so we can write this: \[\Large \begin{gathered} 20 = \frac{1}{8}\left( {\frac{{{v^4}}}{4} + \frac{{{v^2}}}{2}} \right) \hfill \\ \hfill \\ 20 = \frac{1}{8}\left( {\frac{{{v^4} + 2{v^2}}}{4}} \right) \hfill \\ \hfill \\ 640 = {v^4} + 2{v^2} \hfill \\ \hfill \\ {v^4} + 2{v^2} - 640 = 0 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now, we can make this change of variable: w=v^2, so we get: \[\Large {w^2} + 2w - 640 = 0\]

rvc (rvc):

wait 8 X 20 = 160

rvc (rvc):

oh okayyy

OpenStudy (michele_laino):

now, we have to solve with respect to w, and we have to pick the positive solution of w, since we have to remember that w=v^2

OpenStudy (michele_laino):

I got this: \[\Large w = {v^2} = - 1 + \sqrt {640} \cong - 1 + 25.3 = 24.3\]

OpenStudy (michele_laino):

and hence: \[\Large v = \sqrt {24.3} \cong 5.03\] I write the positive solution only for v, since I believe that v is the magnitude of the velocity of our particle

rvc (rvc):

i read every step i understand

OpenStudy (michele_laino):

ok! :)